Mathematics 243 – Algebraic Structures Definitions Quiz 4 – October 6, 2006

Your Name:

A) (5) State the three properties that define equivalence relations.

B) (5) Let R be the relation on Z defined by xRy if (and only if) |x − y| ≤ 2. Is R an equivalence relation? Justify your answer.